# Enhancing universal machine learning potentials with polarizable long-range interactions

### v1

@@ Abstract
Long-range interactions are crucial in determining the behavior of chemical systems in various environments. Accurate predictions of physical and chemical phenomena at the atomic level hinge on accurate modeling of these interactions. Here, we present a framework that substantially enhances the predictive power of machine learning interatomic potentials by incorporating explicit polarizable long-range interactions with an equivariant graph neural network short-range potential. The pretrained universal model, applicable across the entire periodic table, can achieve first-principles accuracy. This versatile model has been further applied to diverse areas of research, including the study of mechanical properties, ionic diffusivity in solid-state electrolytes, ferroelectricity, and interfacial reactions, demonstrating its broad applicability and robustness.
Main
@@
While existing MLIPs with a cut-off of around 5 Å perform well in simulating interactions within localized chemical environments, they may fail to capture long-range phenomena. This hinders their ability to understand and elucidate the behaviors of complex materials18. These behaviors encompass electrostatic interactions, involving forces between charged particles, and dispersion terms, which arise from subtle fluctuations in electron distribution within molecules or atoms. Consequently, this limitation calls for further research and development to enhance the predictive capabilities of MLIPs across a wider range of distances and interactions. Various models have been proposed to address the challenges associated with long-range interactions. The third-generation19, 20 high-dimensional neural network potentials (HDNNPs) consider partial charges to be local properties, which could potentially be insufficient for describing systems with strong nonlocal dependencies in their electronic structures. To overcome this limitation, fourth-generation21, 22 HDNNPs were developed, incorporating the charge equilibration (QEq)23 principles. In fourth-generation HDNNPs, partial charges are directly trained to match reference charges derived from first principles, resulting in qualitatively meaningful outcomes. This new approach has shown early success in addressing the shortcomings of the third generation.
The polarizable charge equilibration (PQEq)24, 25, 26 method proposed by Naserifar et al. enhances the QEq method by incorporating polarization effects27, 28. Unlike the fourth-generation HDNNPs and the conventional QEq scheme, PQEq uses quantum mechanics (QM) interaction energies instead of partial charges as targets for evaluating interatomic potentials. Notably, partial charges obtained from QM calculations are either derived from the direct partitioning of the molecular wave function into atomic contributions or calculated based on the analysis of physical observables. However, due to the incompleteness of basis sets and variances in the partitioning methods, properties calculated based on partial charges may not always be reliable16, 24. Hence, substituting partial charges with the QM electrostatic interaction energies is expected to yield more general and widely applicable results. In previous works, there has been the development of PQEq parameters for 102 elements, derived from experimental data or high-level QM calculations23, 29, and demonstrated its effectiveness in accurately reproducing QM electrostatic interaction energies24.
@@
In this work, we introduce a novel framework that integrates the equivariant neural network potentials with the polarizable long-range electrostatic potentials. When combined with dispersion potentials, the pretrained universal model within this framework achieves accuracy comparable to first principles methods for elements in the periodic table up to Pu, while maintaining a low computational cost. Furthermore, we have applied this universal pretrained model across different areas of applications, including the predictions of mechanical properties of materials, phase transitions and hysteresis behaviors in ferroelectric materials, and interfacial reactions in solid-state batteries, thus leading to significant advancements in the fields of materials science and chemical simulation.
@@ Results
Framework and performance
Figure 1: The architecture of the framework and model performance on bulk modulus. a, The framework integrates short-range and long-range interactions, encompassing nearly the entire periodic table up to plutonium. Atomic number one-hot encoding, radial basis functions, and spherical harmonics are employed to construct the short-range atomic graph. This graph is then processed through an equivariant graph neural network to output the short-range energy. The lower part details the calculation of long-range interaction energy, which includes electrostatic and DFT-D3 dispersion terms. Each atom is represented by the Gaussian-type core and shell, enabling the calculation of electrostatic energy and partial charges through polarizable charge equilibration. The model can predict the forces and stresses by applying automatic differentiation to the total energy. Benchmarks on various calculation methods for the bulk modulus derived from first principles calculations and the model: b, Voigt approach, c, Reuss method, and d, Hill average.
@@
Our framework, which integrates short-range and long-range interactions is depicted in Fig. 1a. For a given chemical system and the boundary condition, our goal is to construct a mapping from atomic coordinates $\bm{r_{i}}$ and atomic types $z_{i}$ to the total potential energy $E_{\rm pot}$. The potential energy is derived from second-order expansion24 with respect to charge fluctuations and augmented by the dispersion term $E_{\rm D3}$,
$\begin{split}E_{\rm pot}&=\sum_{i}\left(E_{0}^{i}(\bm{r_{i}},z_{i})+\chi_{i}^{0}q_{i}+\frac{1}{2}\eta_{i}^{0}q_{i}^{2}+\frac{1}{2}K_{s}^{i}\bm{r_{ic,is}}^{2}\right)\\ &+\sum_{i>j}C_{ik,jl}(\bm{r_{ik,jl}})q_{ik}q_{jl}+E_{\rm D3}.\end{split}$ (1)
@@
The zero-order atomic energy $E_{0}^{i}$ corresponds to the short-range interactions of atom $i$, i.e., the final layer scalar output of the equivariant graph neural networks, while the higher-order self and interatomic Coulomb interactions represent the long-range electrostatic interactions. In the equivariant neural networks12, the initial features are generated using a trainable one-hot embedding that operates on the atomic types. The interatomic distance of atom $i$ and atom $j$, denoted as $|\bm{R_{ij}}|$, is expanded by radial basis functions. Concurrently, the directional component of the interatomic vectors, expressed as $\frac{\bm{R_{ij}}}{|\bm{R_{ij}}|}$, is expanded by spherical harmonics functions. These features are utilized to construct the atomic graph. Then, through the equivariant message passing30 schemes, the atomic features are updated. A multilayer perceptron layer is then connected to derive the scalar output. To account for charge transfer and polarization effects, partial atomic charge $q_{i}$ of atom $i$ is contributed by the sum of the nuclear core charge $q_{ic}$ and shell charge $q_{is}$, both of which assume a Gaussian charge distribution form. The first-order coefficients $\chi^{0}$ are electronegativities, commonly defined as half of the sum of ionization potential (IP) and electron affinities (EA). The second-order coefficients $\eta^{0}$ signifies idempotential or chemical hardness, defined as $IP-EA$. The spring constant $K_{s}^{i}$ denotes the isotropic harmonic connectivity between the shell position $\bm{r_{is}}$ and core position $\bm{r_{ic}}$ of atom $i$. The Gaussian electrostatic energy is given by $C(\bm{r_{ik,jl}})q_{i}kq_{j}l$, where $i$ and $j$ are the atomic indices, and $k$ and $l$ represent the core ($c$) or shell ($s$), respectively. In MD simulations, the partial charges are updated dynamically by minimizing the potential energy equation (Eq. 1) subject to the conservation of the net charge, and to the equality of chemical potentials for all atoms. The detailed derivation of PQEq and the response of the force to the external electric field can be found in the Supplementary Information Part 1. As our goal is to reproduce Density Functional Theory31 (DFT) results, the DFT-D3 van der Waals dispersion energies correction32, 33 terms $E_{\rm D3}$ are employed to account for the dispersion. The forces on atoms $\bm{F_{i}}$ and stress $\sigma_{ab}$ on the cell can be calculated via auto differentiation of potential energy with respect to atomic coordinates and strain $\epsilon_{ab}$,
Figure 2: Li diffusivity in cubic phase Li7La3Zr2O12. a, Crystal structure of $\rm Ia\bar{3}d$ $\rm Li_{7}La_{3}Zr_{2}O_{12}$ and Arrhenius plots depicting the lithium-ion diffusion coefficients across varying temperatures. The dark blue polyhedron signifies La located at the 24(c) site and the light brown polyhedron indicates Zr at the 16(a) site. Li fraction occupies the 24(d) and 96(h) sites. Predicted diffusion coefficients of AIMD34 and our model, complete with error bars, are presented to calculate activation energies. b, 2-ns mean square displacements of lithium-ion in c-LLZO with different temperatures ranging from 800 K to 1800 K in an increment of 200K. The linear dashed gray lines, with a slope of 1, are also plotted.
$\bm{F_{i}}=-\frac{\partial E_{\rm pot}}{\partial\bm{r_{i}}}$ (2)
$\sigma_{ab}=\frac{1}{V_{\rm cell}}\frac{\partial E_{\rm pot}}{\partial\epsilon_{ab}},$ (3)
where $V_{\rm cell}$ denotes the cell volume.
@@
We trained a universal model for all the periodic table elements up to Pu, as described in the Methods section, to evaluate the feasibility of this framework. Our results demonstrate mean absolute errors (MAE) of 18 meV per atom for energies $E$, 0.064 eV/Å for forces $\bm{F}$, and 0.301 GPa for stresses $\bm{S}$ across a test set comprising 155,765 configurations. These results show a modest improvement compared to the previously established pretrained universal model16, suggesting that incorporating long-range interactions enhances the accuracy of model in reproducing the outcomes predicted by DFT. To demonstrate the performance of the model, we conducted benchmarking using properties that were not labeled during the training process. First, we applied the pretrained model to predict the mechanical properties of 10,154 materials from the Materials Project35. Fig. 1b-d illustrate the comparison of the bulk modulus $B$ determined by the pretrained model and DFT. Our pretrained model demonstrated impressive performance, achieving an $R^{2}$ of 0.94 with the Voigt approach36 $B_{\rm V}$ and Hill average method37 $B_{\rm VRH}$, and 0.93 using the Reuss method38 $B_{\rm R}$. These achievements not only highlight the robust performance of the model but also establish a solid foundation for the high-throughput screening of materials with exceptional mechanical properties.
Li diffusion in solid-state electrolytes
Figure 3: The phase transitions and P-E loop of BaTiO3. The temperature dependence of a, lattice constants, and b, local polarizations of unit cells exhibit notable changes during the phase transitions observed from MD simulations on 10×10×10 BaTiO3. c, P-E loop of 20×10×10 tetragonal BaTiO3 at 235 K. The electric field is applied along the x-direction.
@@
Next, we simulate the kinetic transport properties of a solid-state electrolyte to demonstrate the capability of our model to reproduce results from AIMD simulations. Specifically, we investigated lithium-ion diffusivity within the cubic phase of $\rm Li_{7}La_{3}Zr_{2}O_{12}$ (c-LLZO)39, a crystalline superionic conductor known for its remarkable stability as a lithium conductor. Owing to the efficiency of our pretrained model, we are able to conduct MD simulations on the c-LLZO comprising 64 formula units, with a duration of 2 ns for each temperature range from 800 K to 1800K. Fig. 2a presents the Arrhenius plots derived from our model and AIMD34 simulations. The diffusion coefficients were determined from the slope of the logarithmic mean square displacements (MSD) versus logarithmic time within the Fickian regime40, as depicted in Fig. 2b. Compared to previous work using AIMD, our model clearly reproduces the diffusion coefficients and activation energy of c-LLZO. More importantly, with the ability to reach larger model sizes and longer simulation times, our model achieves these results with significantly reduced error. This capability not only delivers comparable results to AIMD simulations but also enables larger-scale simulations that cannot be assessed by AIMD simulations, enhancing the statistical significance of MD studies. Consequently, the framework developed in this work provides more reliable and comprehensive data for diffusion analysis in solid-state electrolytes.
Polarization and ferroelectricity
To verify the ability of our model to handle polarized systems, we performed MD simulations on a typical perovskite ferroelectric material $\rm BaTiO_{3}$. As temperature increases, $\rm BaTiO_{3}$ undergoes a non-trivial phase transition process41. Below 183 K, it adopts a rhombohedral structure with local polarization along the [111] direction. As the temperature rises, it transitions to an orthorhombic phase with polarization along the [110] direction, then to a tetragonal phase with polarization along [100] direction at 278 K. Finally, at 403 K, it converts to a cubic paraelectric phase. Accurately simulating phase changes in ferroelectric substances requires precise potential energy functions that can respond to the small atomic shifts and structural changes, as well as account for the free energy surfaces under the conditions of finite-temperature thermodynamics42. The rhombohedral-orthorhombic-tetragonal-cubic phase sequence of $\rm BaTiO_{3}$ has been extensively studied and reproduced using effective Hamiltonians43, 44, CMD45, 46, and specialized MLIPs models442, 47.

### v2

# A foundation machine learning potential with polarizable long-range interactions for materials modelling

@@ Abstract
Long-range interactions are essential determinants of chemical system behaviour across diverse environments. We present a novel framework that integrates explicit polarizable long-range physics with an equivariant graph neural network potential. It employs a physically motivated polarizable charge equilibration scheme that directly optimizes electrostatic interaction energies rather than partial charges. The foundation model, trained across the periodic table up to Pu, demonstrates strong performance across key materials modelling challenges. It effectively captures long-range interactions that are challenging for traditional message-passing mechanisms and accurately reproduces polarization effects under external electric fields. We have applied the model to mechanical properties, ionic diffusivity in solid-state electrolytes, ferroelectric phase transitions, and reactive dynamics at electrode-electrolyte interfaces, highlighting the model’s capacity to balance accuracy and computational efficiency. Furthermore, we show that as a foundation model, it can be efficiently finetuned to achieve high-level accuracy for specific challenging systems.
Introduction
@@
While existing local MLIPs with a cut-off of around 5 Å perform well in simulating interactions within localized chemical environments, they may fail to capture long-range phenomena. This hinders their ability to understand and elucidate the behaviours of complex materials 18. An approach is to implicitly incorporate the effects of long-range interactions through message-passing mechanisms 19, which can extend the cutoff through layer-wise propagation. However, if certain parts of the system are disconnected on the graph, such as two molecules separated by a distance beyond the cutoff value, the message-passing scheme becomes ineffective 20. A promising direction involves investigating the potential advantages of explicitly including long-range interactions in the model formulation. These behaviours encompass electrostatic interactions, involving forces between charged particles, and dispersion terms, which arise from dynamic fluctuations in electron distribution within molecules or atoms.
Various explicit models have been proposed to address the challenges associated with long-range interactions by including electrostatic interactions. In these methods, charges are usually obtained through charge equilibration (QEq) 21 principles, including the Charge Equilibration via Neural Network Technique (CENT) method 22, Fourth-generation High-Dimensional Neural Network Potential (4G-HDNNP) 23, 24. Generally, electrostatic parameters were trained to minimise the difference between QEq charges and reference partial charges derived from quantum mechanics (QM) calculations. Notably, partial charges obtained from QM calculations are either derived from the direct partitioning of the molecular wave function into atomic contributions or calculated based on the analysis of physical observables. However, due to the incompleteness of basis sets and variances in the partitioning methods, properties calculated based on partial charges may not always be reliable 25. The ambiguity in DFT-assigned charges suggests that directly learning of them may be inessential for constructing accurate interatomic potentials. On the contrary, the polarizable charge equilibration (PQEq) 25, 26, 27 method proposed by Naserifar et al. enhances the QEq method by using QM electrostatic interaction energies instead of partial charges as targets for evaluating interatomic potentials. Additionally, PQEq has proven to be effective in capturing polarization effects by explicitly introducing core-shell displacement. Another way to solve the partition-dependent issue is to take dipole moment into the loss function, as done by SpookyNet 28. However, it still uses point-like charges, making it behaves like QEq-based methods when dealing with polarization effects.
@@
In this work, we introduce a novel framework that integrates the equivariant message-passing neural network potentials with the explicit polarizable long-range potential. We initially conducted benchmark tests on datasets containing systems with various charge states proposed by references 24, 29, showcasing the efficacy of our framework in the handling both long-range interactions and varying charge environments. Building on these results, we extended the framework to develop a foundation model encompassing the periodic table up to Pu, while maintaining computational efficiency. We have applied this foundation model across different areas of applications, including the predictions of mechanical properties of materials, phase transitions in ferroelectric materials, and reactive molecular dynamics of solid-state batteries interface, thus leading to significant advancements in the fields of materials science and chemical simulations. The model also demonstrates transferability, accurately predicting interactions between clusters in various charge states, and molecular polarizability. Furthermore, we show that the foundation model can be efficiently finetuned to achieve ab initio accuracy for specific challenging systems.
@@ Results
Theoretical framework
Figure 1: a, Overview of the framework. The framework takes atomic numbers ($Z_{i}$) and coordinates ($\bm{r}_{i}$) as inputs and integrates neural network block and explicit polarizable long-range interactions block. For the neural network block, atomic numbers are converted into feature vectors via one-hot encoding, while the coordinates are transformed into a representation of the local environment using radial basis functions and spherical harmonics to capture the distance ($|\bm{R}_{ij}|$) and directional ($\bm{R}_{ij}/|\bm{R}_{ij}|$) information. These embeddings are processed through an equivariant graph neural network to output the scalar MLIP energy $E_{0}$. Explicit, physically-grounded calculations are performed for long-range effects. An electrostatic energy component ($E_{\rm PQEq}$) is calculated using the polarizable charge equilibration (PQEq) method, and a dispersion energy component ($E_{\rm D3}$) is calculated using the DFT-D3 method. The total energy is the sum of these components. Gradients of the energy yield forces and stress on the system. The long-range block also provides partial charges as an auxiliary output. b, A flowchart detailing the iterative self-consistent procedure for the PQEq method. Starting with atomic core positions ($\bm{r}_{ic}$), atomic numbers ($Z_{i}$), and predetermined atomic parameters (electronegativities $\chi_{i}^{0}$, chemical hardness $\eta_{i}^{0}$, and spring constant $K_{s}^{i}$), the method uses a core-shell Gaussian model to build and solve a system of linear equations. This loop is repeated until the partial charges ($q_{i}$) and shell positions ($\bm{r}_{is}$) converge, yielding the final electrostatic energy ($E_{\rm PQEq}$) and the converged partial charges. c, Partition of core-shell Gaussian charge model used in long-range electrostatics. The partial charge ($q_{i}$) of each atom is represented as sum of the core ($q_{ic}$) and shell ($q_{is}$) charge. The harmonic spring constant, $K_{s}^{i}$, couples the core and shell.
@@
Our framework, which integrates neural network potential and polarizable long-range interactions, is depicted in Fig. 1. For a given chemical system and boundary conditions, our goal is to construct a mapping from atomic coordinates $\bm{r}_{i}$ and atomic types $Z_{i}$ to the total potential energy $E_{\rm pot}$. The potential energy is expressed as the sum of the second-order expansion 25 with respect to charge fluctuations and the density functional theory 30 D3 (DFT-D3) van der Waals dispersion energies correction 31 term $E_{\rm D3}$,
$\begin{split}E_{\rm pot}&=\sum_{i}\left(E_{0}^{i}(\bm{r}_{i},z_{i})+\chi_{i}^{0}q_{i}+\frac{1}{2}\eta_{i}^{0}q_{i}^{2}+\frac{1}{2}K_{s}^{i}\bm{r}_{ic,is}^{2}\right)\\ &+\sum_{i>j}C_{ik,jl}(\bm{r}_{ik,jl})q_{ik}q_{jl}+E_{\rm D3}\\ &=\sum_{i}E_{0}^{i}+E_{\rm PQEq}+E_{\rm D3}.\end{split}$ (1)
@@
The zeroth-order atomic energy $E_{0}^{i}$ corresponds to the final layer scalar output of the equivariant graph neural networks, while the higher-order self- and interatomic Coulomb interactions represent the polarizable long-range electrostatic interactions $E_{\rm PQEq}$. To account for charge transfer and polarization effects, partial atomic charge $q_{i}$ of atom $i$ is the sum of the nuclear core charge $q_{ic}$ and shell charge $q_{is}$, both of which assume a Gaussian charge distribution form. The first-order coefficients $\chi^{0}$ are electronegativities, commonly defined as half of the sum of ionization potential (IP) and electron affinities (EA). The second-order coefficients $\eta^{0}$ signifies idempotential or chemical hardness, defined as IP-EA. The spring constant $K_{s}^{i}$ denotes the isotropic harmonic connectivity between the shell position $r_{is}$ and core position $r_{ic}$ (equal to $r_{i}$) of atom $i$. The Gaussian electrostatic energy is given by $C(\bm{r}_{ik,jl})q_{ik}q_{jl}$, where $i$ and $j$ are the atomic indices, and $k$ and $l$ represent the core ($c$) or shell ($s$), respectively. The displacement vector is given by $\bm{r}_{ik,jl}=\bm{r}_{ik}-\bm{r}_{jl}$. The detailed derivation of the response to the external electric field can be found in the Supplementary Information. In previous works, PQEq parameters for 102 elements have been derived from experimental data or high-level QM calculations 25, and have accurately reproduced QM electrostatic interaction energies. In the equivariant neural networks 12, the initial features are generated using a trainable one-hot embedding that operates on the atomic types. The interatomic distance of atom $i$ and atom $j$, denoted as $|\bm{R}_{ij}|$, is expanded by radial basis functions. Concurrently, the directional component of the interatomic vectors, expressed as $\bm{R}_{ij}/|\bm{R}_{ij}|$, is expanded by spherical harmonics functions. These features are utilized to construct the atomic graph. Then, through the equivariant message passing 19 schemes, the atomic features are updated. A multilayer perceptron layer is then connected to derive the scalar output. It should be noted that the neural network potential component is specifically designed to capture energy contributions distinct from electrostatic interactions, thereby ensuring no overlap in energy accounting between different components. Subject to the conservation of the net charge and the equality of chemical potentials for all atoms, the linear equations can be used to updated partial charges ($q_{i}$) and shell positions ($\bm{r}_{is}$). The partial charges can be obtained self-consistently in each MD step. The forces on atoms and virial stress on the cell can be calculated via automatic differentiation after partial charges are determined as detailed in Supplementary Information.
Validation on diverse charge-state datasets
[…]
Table 1: The comparison of root means square test error metrics on the different charge-state datasets.
Dataset 4G-HDNNP Maruf’s NequIP This work
$\rm{C_{10}H_{2}/C_{10}H_{3}^{+}}$ Energy (meV/atom) 1.194 1.33 0.44
Force (eV/Å) 0.078 0.071 0.023
[…]
Force (eV/Å) 0.033 N.A. 0.005
[…]
Force (eV/Å) 0.032 2.145 0.028
Au2-MgO Energy (meV/atom) 0.22 1.03 0.14
Force (eV/Å) 0.066 0.034 0.021
BTA-Cu Energy (meV/atom) N.A. 0.48 0.30
Force (eV/Å) N.A. 0.008 0.004
BTA (H2O)-Cu Energy (meV/atom) N.A. 0.71 0.19
Force (eV/Å) N.A. 0.010 0.005
For the carbon chains, the introduction of additional H+ causes some C atoms to exhibit opposite charge states, as shown in Fig. S1 a. Although we did not fit partial charges, our model qualitatively captures the distribution of DFT partial charges, as shown also in Fig. S1 a. In the linear C10H2 molecule, our model demonstrates symmetric charge distribution around the centre which shows agreement with DFT. After explicitly incorporating physical long-range interactions, our framework achieved additional improvements in error metrics compared to NequIP and 4G-HDNNP. For positively charged NaCl clusters, the predicted energy and forces by our model are still more accurate compared to 4G-HDNNP, as well as the potential energy surface shown in Fig. S1 b. For Au2 adsorption on MgO surfaces and benzotriazole interactions with Cu-(111) surfaces in dry and aqueous environments, our model outperforms NequIP significantly. Additionally, our model correctly predicted the preferential adsorption geometries for both doped and undoped MgO, as listed in Table S1. As for the charged Ag clusters, due to their small system size and lacking long-range effects, the entire structure falls within the typical cutoff radius used by the MLIP. The energy dependence on the overall charge state of clusters leads to degeneracy between atomic structures and potential energy surfaces, resulting in poor performance of charge-independent NequIP model. We acknowledge that for such small, highly charged systems, 4G-HDNNP approach to learning charge equilibration parameters is more convenient. Nonetheless, we also achieved reduced force error with the fixed physical parameters compared to 4G-HDNNP. Although PQEq partial charges are not directly comparable to DFT Hirshfeld ones, we compare them to show their relevance. Varying degrees of agreement across different chemical systems are observed as shown in Table S2. Our comprehensive evaluation demonstrates that explicit incorporation of physical long-range interactions significantly enhances the performance of MLIPs across diverse charge-state systems. Notably, our proposed framework outperforms 4G-HDNNP in most cases for energies and forces without fitting partial charges. This advantage likely stems from a more physically meaningful partition of the total energy.
Foundation model benchmark
Figure 2: a, The interaction energies for Na-Na (green) and Na-Cl (purple) dimers predicted by our model, w/o-lr model, and reference calculations using DFT. b, Evaluation of polarizable interactions in water molecules. The water molecule model oriented in the yz-plane, with oxygen and hydrogen atoms shown in red and white, respectively. The energy response curves as a function of external electric field ($\bm{\epsilon}$) strength applied along the x-axis. Results are compared among our model with polarizable long-range interactions (blue), the QEq-based model (yellow), and the reference DFT calculations (green) serving as the reference values. c-e, Comparison of bulk modulus from first principles calculations and our model: c, Voigt approach, d, Reuss method, and c, Hill average.
We trained a foundation model for all the periodic table elements up to Pu using MPtrj dataset 16, as described in the Methods section, to evaluate the feasibility of this framework. We also trained a model maintaining the same MLIPs framework and dataset but without the polarizable long-range interactions (w/o-lr model). The former one demonstrated mean absolute errors (MAE) of 18 meV per atom for energies, 0.065 eV/Å for forces, and 0.301 GPa for stresses across the test set, outperforming the w/o-lr model, as shown in Table S4. These proved the benefit of incorporating long-range interactions in the enhancement of prediction accuracy.
Message-passing mechanisms will fail when graph connections break in the system. For instance, in dimers, when the interatomic distance exceeds the graph cut-off, the atoms no longer interact with each other. We analysed interaction energies for Na-Na and Na-Cl dimers across interatomic distances ranging from 1.8 Å to 12.0 Å. As shown in Fig. 2 a, both models (our model and w/o-lr model) trained on MPtrj dataset accurately reproduce the equilibrium bond lengths of the two dimers. However, the w/o-lr model failed to differentiate interaction energies beyond the MLIPs cutoff distance, i.e., 5.0 Å, while our model successfully captured the DFT potential energy surface throughout the entire distance range, demonstrating its capability in describing extended ionic interactions. Moreover, the message-passing mechanism exhibits deficiencies when simulating layered materials. Taking lithium iron phosphate (LiFePO4) as a case study, we observed that the w/o-lr model predicted an irreversible phase transition as shown in Fig. S3. In contrast, our model correctly reproduced the experimentally 32 and AIMD 33 observed thermally stable olivine structure. These results collectively demonstrate that the explicit long-range interactions are crucial for accurately describing both ionic systems and layered materials, particularly in predicting structural stability and thermal behaviour.
Polarization plays a fundamental role in determining molecular responses to external electric fields. Conventional QEq-based models have limitations in their response to an external electric field applied orthogonally to the molecular dipole. We validated this capability through a benchmark study examining a water molecule’s response to external electrostatic fields 34. The water molecule was fixed in the $yz$-plane, and we applied varying external electric fields along the $x$-axis to derive energy curves, with the results presented in Fig. 2 b. The model with polarizable long-range interactions demonstrates a trend consistent with the reference values obtained from DFT. Due to the external electrostatic field direction consistently being orthogonal to the molecule, the electrostatic energy in the QEq-based model (like 4G-HDNNP) is 0 and is completely independent of field strength ($i.e.,\bm{\mu}\cdot\bm{\epsilon}=0$, where $\bm{\mu}$, $\bm{\epsilon}$ are dipole moment and electric field, respectively). This is because the QEq model treats atoms in the molecule as point charges or Gaussian charges, without distinguishing between core and shell components. We also compared the response of O atom partial charges across DFT Hirshfeld calculations, our model, and the conventional QEq model as shown in Fig. S4. The agreement between PQEq model and DFT is observed, while QEq model completely lacks response. This comparative analysis establishes that the explicit polarizable interactions are crucial for modelling molecular systems subject to external electric fields.
Besides the failure of response in static conditions, the QEq method may exhibit non-physical behaviour in dynamics 35. As illustrated in Fig. S5, we conducted additional investigations of two water molecules under an external electric field of 0.25 V/Å. In charge distribution models, it is essential to allow systems to form an internal electric field opposing the external field. The QEq model assumes the system behaves as a perfect conductor without penalizing charge transfer as a function of distance. This limitation of the QEq method leads to non-physical charge transfer between molecules separated by large distances 35. In contrast, within PQEq model, the individual water molecules can polarize, generating internal electric fields that counteract the external field, thereby capturing more realistic electrostatic responses in molecular systems. Under QEq scheme, the water molecules accumulate non-zero net charges and migrate to the top and bottom of the simulation box. In contrast, our model maintains charge neutrality of the water molecules, which remain stationary despite the applied electric field, aligning with DFT results.
@@
We conducted benchmarking using properties that were not labeled during the training process. First, we applied the foundation model to predict the mechanical properties of 10,154 materials from the Materials Project36. Fig. 2 c-e illustrate the comparison of the bulk modulus $B$ determined by the foundation model and DFT. Our foundation model demonstrated impressive performance, achieving an $R^{2}$ of 0.94 with the Voigt approach37 $B_{\rm V}$ and Hill average method38 $B_{\rm VRH}$, and 0.93 using the Reuss method39 $B_{\rm R}$. In comparison, as shown in Fig. S2, the w/o-lr model achieved lower $R^{2}$ values of 0.89, 0.86, and 0.88 for $B_{\rm V}$, $B_{\rm R}$, and $B_{\rm VRH}$, respectively. These not only highlight the robust performance of the model but also establish a solid foundation for the high-throughput screening of materials with exceptional mechanical properties.
Transferability of the foundation model
In terms of foundation model transferability, evaluations were conducted across diverse systems. The model accurately reproduces the potential energy surfaces and successfully distinguishes between neutral and ionic states in OH-OH systems, capturing the long-range Coulombic repulsion of ionic state as shown in Fig. S8. Further validation using a periodic water system demonstrates the model’s capability to predict polarization effects under varying electric fields, achieving comparable performance to specialized models as depicted in Fig. S9. Additionally, as shown in Fig. S10, when tested on 7,211 molecules from the QM-7b dataset 40, the model shows remarkable accuracy in predicting molecular polarizability with a mean absolute error of 4.57 a.u., highlighting its robust transferability in capturing polarization.
Computational efficiency
Despite the incorporation of additional long-range interaction calculations, our foundation model maintains high computational efficiency. Performance benchmarks were conducted on a single NVIDIA H100 GPU to quantitatively assess the computational cost. As illustrated in Fig. S6, for a periodic system comprising 2160 atoms, our model requires approximately 0.07 s per molecular dynamics time step, representing a significant improvement over conventional universal MLIPs which require approximately 0.91 s. This computational advantage extends to larger systems, as demonstrated by simulations of a 24000-atom system, where our model maintains efficiency with only 1.49 s per time step. Such computational performance enables nanosecond-scale molecular dynamics simulations of systems containing tens of thousands of atoms, making it practical for more realistic modelling of materials.
Materials modelling applications
Figure 3: a, Crystal structure of $\rm{Ia\bar{3}d}$-Li7La3Zr2O12 and Arrhenius plots depicting the lithium-ion diffusion coefficients across varying temperatures. The dark blue polyhedron signifies La located at the 24(c) site and the light brown polyhedron indicates Zr at the 16(a) site. Li fraction occupies the 24(d) and 96(h) sites. Predicted diffusion coefficients of AIMD 41, our model, and w/o-lr model complete with error bars, are presented to calculate activation energies. b, 2-ns mean square displacements using our model of lithium-ion in c-LLZO with different temperatures ranging from 800 K to 1800 K in an increment of 200K. The linear dashed grey lines, with a slope of 1, are also plotted.
@@
To demonstrate the capability of our model to reproduce results from AIMD simulations, we simulate the kinetic transport properties of a solid-state electrolyte. We investigated lithium-ion diffusivity within the cubic phase of $\rm Li_{7}La_{3}Zr_{2}O_{12}$ (c-LLZO)42, a crystalline superionic conductor known for its remarkable stability as a lithium conductor. We conduct MD simulations on the c-LLZO comprising 64 formula units, with a duration of 2 ns for each temperature range from 800 K to 1800K. Fig. 3 a presents the Arrhenius plots derived from our model and AIMD41 simulations. The diffusion coefficients were determined from the slope of the logarithmic mean square displacements (MSD) versus logarithmic time within the Fickian regime43, as depicted in Fig. 3 b. Compared to previous work using AIMD simulations, our model clearly reproduces the diffusion coefficients and activation energy. The performance surpasses that of the w/o-lr model which lacks explicit long-range interactions. With the ability to reach larger model sizes and longer simulation times, our model achieves these results with significantly reduced uncertainties. This capability not only delivers comparable results to AIMD simulations but also enables larger-scale simulations that cannot be assessed by AIMD, enhancing the statistical significance of MD studies. Consequently, the framework developed in this work provides more reliable and comprehensive data for diffusion analysis in solid-state electrolytes.